# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

### v4

@@ 5.2 Quantitative Comparison
We evaluate the stylized videos generated by different methods across four aspects, in order to comprehensively assess these methods’ performance. (1) Overall transfer performance. Drawing inspiration from the methodologies outlined in [17, 30], we assess the overall transfer performance using ArtFID [48], a comprehensive metric that combines LPIPS [49] for gauging content fidelity and FID [50] for assessing style richness. (2) Local style preservation. To evaluate the preservation of local style, we utilize SSIM [51] metric for the texture preservation, while CLIP-I score [52] is utilized to assess the preservation of semantic content. During the evaluation, the background is neutralized to ensure accuracy. (3) Temporal consistency. In line with the approaches detailed in [30, 11], we measure the similarity between consecutive frames using the CLIP-F score [52], which serves as a measure of inter-frame coherence in the edited video. (4) Inference efficiency. To assess the inference efficiency, we calculate the aggregate time during inference, thereby assessing their performance in terms of computational speed. Furthermore, we employ six established methods as baselines and benchmark them against our UniVST on two datasets: DAVTG-WikiArt and DAVTG-Laion. Quantitative comparisons are comprehensively shown in Table I and Table II.
The results demonstrate that our UniVST outperforms all competing methods across three performance metrics. On both the DAVTG-WikiArt and DAVTG-Laion datasets, UniVST consistently surpasses existing methods in local style preservation and temporal consistency, showcasing its superiority in localized video style transfer tasks over existing methods. Regarding overall transfer performance, UniVST achieves the best results on the DAVTG-Laion dataset and secures the second spot on DAVTG-WikiArt. Localized transfer tasks might lead to a marginally higher FID score, slightly affecting overall performance compared to other methods. Nevertheless, these findings underscore UniVST’s proficiency in effectively balancing content fidelity and style richness.
@@
TABLE III: Ablation study on the DAVTG-WikiArt dataset. Bold values indicate the best performance.
Model Variants Overall Transfer Performance Local Style Preservation Temporal Consistency
@@
W/o all components (baseline) 45.676 31.380 0.411 0.979 0.976 0.981
W/o localized latent blending 38.014 26.050 0.405 0.966 0.970 0.979
W/o AdaIN-guided latent-shift 37.649 26.909 0.349 0.985 0.990 0.980
W/o AdaIN-guided attention-shift 39.589 27.506 0.388 0.985 0.988 0.983
W/o sliding-window consistent smoothing 38.626 26.907 0.384 0.982 0.989 0.970
W/ all components (ours) 37.152 26.134 0.369 0.986 0.991 0.980
5.3 Qualitative Comparison
For qualitative evaluation, we have juxtaposed the editing results with those of six baseline methods, as illustrated in Fig. 6. Diffutoon [12], which is renowned for its expertise in cartoon coloring, exhibits limited responsiveness to a variety of image styles. This results in a final edited video that is notably monotonous, and heavily inclined towards a cartoonish anime aesthetic. StyleCrafter [11], tailored for style-based video generation, demonstrates a high sensitivity to stylistic cues. However, it encounters difficulties in preserving content fidelity, leading to substantial distortion of background details and character features during the transfer process. BIVDiff [10] falls short in both content fidelity and temporal consistency, similarly altering background information and character traits during the transfer. AnyV2V [13] excels in maintaining temporal consistency but faces challenges with transfer tasks involving a central object. It often over-transfers the subject’s style, resulting in a loss of its distinctive attributes. As for the traditional non-diffusion methods, EFDM [20] and CAST [19], they degrade fine details and induce color artifacts during the style transfer process. On the contrary, our UniVST exhibits a remarkable ability to balance content fidelity and style richness, achieving a harmonious integration of both. It ensures that the key structural and motion details of the original videos are retained, while preserving the primary style of the object with great precision.
TABLE IV: Ablation study on the DAVTG-Laion dataset. Bold values indicate the best performance.
Model Variants Overall Transfer Performance Local Style Preservation Temporal Consistency
@@ 5.3 Qualitative Comparison
W/o all components (baseline) 38.981 26.594 0.413 0.979 0.976 0.981
W/o localized latent blending 30.848 20.974 0.404 0.961 0.967 0.979
W/o AdaIN-guided latent-shift 31.001 21.766 0.362 0.985 0.989 0.980
W/o AdaIN-guided attention-shift 32.534 22.623 0.377 0.985 0.989 0.981
W/o sliding-window consistent smoothing 30.834 21.249 0.386 0.981 0.988 0.971
W/ all components (ours) 30.636 21.146 0.383 0.986 0.990 0.981
Figure 7: Visual results for ablations. Various components in our framework have all played important roles. Best view with zooming-in.
5.4 Ablation Studies
In this section, we first validate the effectiveness of each core component in our proposed UniVST. Then, we conduct an in-depth analysis of hyper-parameter settings for the three modules: point-matching mask propagation, AdaIN-guided localized video stylization, and sliding-window consistency smoothing, individually.
TABLE V: Accuracy and efficiency of mask propagation under different value of $n$ on DAVIS-Laion dataset, Bold fonts represent final configuration.
[…]
n=1 0.5321 0.6505 2.365 (Base)
@@ 5.4 Ablation Studies
Figure 8: Quantitative results of mask propagation with different sampling ratio $r$ on DAVIS [44] dataset. Best view with zooming-in.
TABLE VI: Accuracy and efficiency of mask propagation with different values of $k$ on DAVIS-Laion dataset. Bold fonts represent final configuration.
@@
5.4.1 Ablation Study on Core Components
We have developed six distinct variants to assess the contribution of its individual components: (1) UniVST without all components (baseline), (2) UniVST without localized latent blending, (3) UniVST without AdaIN-guided latent-shift, (4) UniVST without AdaIN-guided attention-shift, (5) UniVST without sliding-window consistent smoothing and (6) UniVST with all components (ours). The quantitative outcomes are presented in Table III and Table IV, while the qualitative results are displayed in Fig. 7. Our mask-guided strategy effectively preserves the foreground style during the stylization process, demonstrating its efficacy quantitatively and qualitatively, without the need for additional models like SAM [35]. The AdaIN-guided latent-shift and attention-shift strategies are essential for style transfer. Notably, omitting the attention-shift strategy leads to incomplete transfer and blurred local details, as shown in Fig. 7. Finally, our sliding-window consistency smoothing scheme enhances temporal consistency and effectively reduces video artifacts. In the following section, we conduct an in-depth study of the three proposed modules on the DAVIS-Laion dataset to further substantiate their individual contributions, respectively.
@@ 5.4.1 Ablation Study on Core Components
Timestep Intervals $\leq 0.6T$ $\leq 0.5T$ $\leq 0.4T$ $\leq 0.3T$ $\leq 0.2T$ $\geq\mathbf{0.4T}$ $\geq 0.5T$ $\geq 0.6T$ $\geq 0.7T$ $\geq 0.8T$
@@
Timestep Intervals $\mathbf{0.1T}\sim\mathbf{0.2T}$ $0.2T\sim 0.3T$ $0.4T\sim 0.5T$ $0.5T\sim 0.6T$ $0.7T\sim 0.8T$ $0.8T\sim 0.9T$
@@
TABLE IX: Comparison of our method with other training-free stylization methods on DAVIS-Laion dataset. Bold values indicate final configuration.
Method K-V Replacement K-V AdaIN Ours ($\beta_{t}\uparrow$) Ours ($\beta_{t}\downarrow$)
[…]
5.4.2 Analysis of Point-Matching Mask Propagation
We first investigate the effect of down-sampling rates on propagation accuracy and inference time using the DAVIS dataset [44]. Detailed results are presented in Fig. 8. As the down-sampling rate $r$ decreases, the efficiency of our mask propagation improves significantly, while maintaining the accuracy of the final propagation results. Next, we also examine the impact of anchor frame selection on propagation accuracy and efficiency. As shown in Table V, increasing the number of anchor frames improves propagation accuracy but also leads to longer propagation times. Moreover, we observe that as the number of frames increases, the accuracy gains diminish. After considering these trade-offs, we choose the first frame and the previous $n=9$ frames as the final hyper-parameter configuration. We also compare the results for different values of the hyper-parameter $k$ in $k$-NN [37]. As shown in Table VI, the highest propagation accuracy is consistently achieved when $k=15$, without sacrificing the overall propagation efficiency.
We present qualitative examples of our mask propagation results in Fig. 9. In the short-term propagation setting, we evaluate scenarios involving fast-moving foreground objects, deformations, and occlusions. Despite these complexities, our method demonstrates precise object tracking and robust mask alignment, consistently preserving accuracy across diverse conditions. For long-term propagation, our approach continues to deliver strong performance. The integration of anchor frames enables stable and accurate mask propagation over extended temporal distances, mitigating typical degradation issues associated with long sequences. However, certain limitations persist, which will be examined in Sec. 5.5.
Figure 10: Visual results of latent-shift with different timestep intervals on DAVIS-Laion dataset. Best view with zooming-in.
Figure 11: Quantitative results of sliding-window consistent smoothing with varying timestep intervals. Best view with zooming-in.
5.4.3 Analysis of AdaIN-Guided Localized Video Stylization
We investigate the optimal time interval for this method in Table VII. It means that applying AdaIN-guided attention-shift in the early and middle stages of denoising can yield better results. Furthermore, we explore the optimal timestep intervals for AdaIN-guided latent-shift, with quantitative results provided in Table VIII. Applying AdaIN-guided latent-shift in the late stages yields better results. In addition, we present visual results of these ablation experiments, as shown in Fig. 10 and Fig. 12. For AdaIN-guided latent-shift, applying it in the early stages of denoising can cause issues such as content blurriness and semantic loss during the transfer process. In contrast, AdaIN-guided attention-shift produces better visual results when applied in the early stages of denoising than in the later stages.
Figure 12: Visual results of attention-shift with different timestep intervals on DAVIS-Laion dataset. Best view with zooming-in.
Figure 13: Visual results of sliding-window consistent smoothing with different window sizes on DAVIS-Laion dataset. Best view with zooming-in.
Finally, to further investigate the effectiveness of our proposed AdaIN-guided attention-shift mechanism and the impact of the monotonicity of the time coefficient $\beta_{t}$, we compare the performance of four configurations: K-V replacement [17], K-V AdaIN, our method with increasing time coefficients, and our method with decreasing time coefficients. As shown in Table IX, using decreasing time coefficient enables the AdaIN-guided attention-shift to outperform other training-free methods, such as K-V replacement, K-V AdaIN, or attention-shift with increasing time coefficients.
@@ 5.5 Addtional Analysis
Bottleneck of Inference Time. In Table I and Table II, our proposed UniVST exhibits the longest inference time among all compared methods, revealing a key computational bottleneck. To better understand this limitation, we perform a detailed breakdown of the inference time distribution across the core modules of UniVST, as illustrated in Fig. 14. These modules include: DDIM Inversion, Point-Matching Mask Propagation, AdaIN-Guided Localized Video Stylization, and Sliding-Window Consistent Smoothing. Our analysis indicates that the primary source of inference overhead arises from the sliding-window consistent smoothing module. This component relies on dynamic optical flow estimation during the denoising phase, which improves temporal consistency but also adds significant computational cost. In contrast, most existing text-to-video (T2V) frameworks adopt temporal attention mechanisms, which are generally more efficient though potentially less robust in handling complex motion. To address this trade-off, a promising alternative is to replace the smoothing module with a 3D VAE decoder that directly reconstructs temporally coherent video from the denoised latent representation, thereby maintaining temporal consistency without incurring additional inference time.
Constraints in Mask Propagation. Although our point-matching mask propagation strategy enables zero-shot object localization, as demonstrated in Fig. 9, it presents several limitations: 1. Challenges in Multi-Object Scenarios. In scenes containing multiple similar objects (e.g., two individuals), the semantic similarity between instances can lead to confusion during propagation. This often results in incorrect mask associations, particularly as the propagation distance increases, causing error accumulation over time. While our current evaluation focuses on diverse yet single-object scenes—where such issues have minimal impact—extending the method to robustly handle multi-object scenarios remains an important direction for future work. 2. Limitations in Fine-Grained Mask Propagation. Although the propagation achieves an average Dice score of approximately $0.87$, it occasionally fails to capture fine object boundaries, leading to incomplete or imprecise tracking. As shown in Fig. 9, while the method effectively handles fast motion and occlusions, it struggles to retain detailed edge information. Nonetheless, our localized latent blending strategy, which performs dynamic region-wise fusion during the denoising phase, effectively mitigates this limitation by reducing reliance on high-precision masks. The mechanism can still preserve the object’s stylization quality even without guidance from high-precision masks.
@@
Figure 16: Qualitative comparisons of our proposed UniVST with some existing commercial models. The first row presents the original video, reference image, and mask, followed by the results of Gen-1 [53] in the second row, Gen-3 in the third row, and our proposed UniVST in the final row.
Figure 17: Multi-step video style transfer: first, the background is stylized, and then the foreground is stylized based on the transformed background.

### v5

@@ 5.2 Quantitative Comparison
To comprehensively assess the performance of different methods, we evaluate their stylized videos across four aspects: (1) Overall transfer performance. Drawing inspiration from the methodologies outlined in [17, 30], we assess the overall transfer performance using ArtFID [52], a comprehensive metric that combines LPIPS [53] for gauging content fidelity and FID [54] for assessing style richness. (2) Local style preservation. To evaluate the preservation of local style, we utilize SSIM [55] metric for the texture preservation, while CLIP-I score [56] is utilized to assess the preservation of semantic content. During the evaluation, the background is neutralized to ensure accuracy. (3) Temporal consistency. In line with the approaches detailed in [30, 11], we measure the similarity between consecutive frames using the CLIP-F score [56], which serves as a measure of inter-frame coherence in the edited video. (4) Inference efficiency. To assess the inference efficiency, we calculate the aggregate time during inference, thereby assessing their performance in terms of computational speed. (5) Overall quality. Following the methodology introduced by VBench [57], we comprehensively evaluate stylized videos in terms of subject consistency, background consistency, motion smoothness, dynamic degree, aesthetic quality, imaging quality, etc., and report the averaged score to measure their overall quality. Furthermore, we employ six established methods as baselines and benchmark them against our UniVST on two datasets: DAVTG-WikiArt and DAVTG-Laion. Quantitative comparisons are comprehensively shown in Table I.
The results demonstrate that our UniVST outperforms all competing methods across three performance metrics. On both the DAVTG-WikiArt and DAVTG-Laion datasets, UniVST consistently surpasses existing methods in local style preservation and temporal consistency, showcasing its superiority in localized video style transfer tasks over existing methods. Regarding overall transfer performance, UniVST achieves the best results on the DAVTG-Laion dataset and secures the second spot on DAVTG-WikiArt. Although localized style transfer tasks may lead to a marginally higher FID score, slightly affecting overall performance compared to other methods, our UniVST still achieves a favorable balance between content fidelity and style richness. Moreover, VBench evaluation further highlights the advantage of our UniVST, where it achieves the highest average score, demonstrating superior performance across multiple perceptual dimensions, including consistency, aesthetics, smoothness, artistic quality, and others.
Figure 7: Additional visual results of UniVST on different backbones: SD-v1.5 [32], SD-v2.1 [32], SD-v3.0 [43], SD-v3.5 [43] and Animatediff [44].
TABLE II: Additional quantitative evaluation results of UniVST on different backbones on the DAVTG-Laion dataset.
Metrics SD-v1.5 SD-v2.1 SD-v3.0 SD-v3.5 Animatediff-v2
[…]
@@
To further evaluate the generalization capability of UniVST, we conduct additional quantitative experiments on the DAVTG-Laion dataset using four backbones: SD-v2.1 [32], SD-v3.0 [43], SD-v3.5 [43], and Animatediff-v2 [44]. The corresponding results are shown in Table II. On SD-v2.1 and Animatediff-v2, UniVST achieves ArtFID, SSIM, CLIP-I, and VBench scores comparable to those on SD-v1.5, indicating stable transfer performance, local style preservation, and overall quality. This highlights the strong generalization capability of UniVST. When applied to the advanced rectified-flow models SD-v3.0 and SD-v3.5, UniVST surpasses its performance on SD-v1.5 across all metrics, demonstrating its scalability with increasingly powerful backbones.
5.3 Qualitative Comparison
For qualitative evaluation, we have juxtaposed the stylized videos with those of six baseline methods, as illustrated in Fig. 6. Diffutoon [12], which is renowned for its expertise in cartoon coloring, exhibits limited responsiveness to a variety of image styles. This results in a final edited video that is notably monotonous, and heavily inclined towards a cartoonish anime aesthetic. StyleCrafter [11], tailored for style-based video generation, demonstrates a high sensitivity to stylistic cues. However, it encounters difficulties in preserving content fidelity, leading to substantial distortion of background details and character features during the transfer process. BIVDiff [10] falls short in both content fidelity and temporal consistency, similarly altering background information and character traits during the transfer. AnyV2V [13] excels in maintaining temporal consistency but faces challenges with transfer tasks involving a central object. It often over-transfers the subject’s style, resulting in a loss of its distinctive attributes. As for the traditional non-diffusion methods, EFDM [20] and CAST [19], they degrade fine details and induce color artifacts during the style transfer process. On the contrary, our UniVST exhibits a remarkable ability to balance content fidelity and style richness, achieving a harmonious integration of both. It ensures that the key structural and motion details of the original videos are retained, while preserving the primary style of the object with great precision. Following MimicMotion [9], we use the slice to further visualize the temporal consistency of stylized videos across different methods, with the results shown in Fig. 6. The fluctuation trend of the slice obtained with our UniVST is most similar to that of the original video, demonstrating that UniVST achieves the smallest temporal consistency loss during the style transfer process.
TABLE III: Ablation studies across different datasets. Bold indicate the best performance.
Model Variants Overall Transfer Performance Local Style Preservation Consistency Overall Quality
@@ 5.3 Qualitative Comparison
DAVTG-WikiArt
W/o all components (baseline) 45.676 31.380 0.411 0.979 0.976 0.981 71.423
W/o localized latent blending 38.014 26.050 0.405 0.966 0.970 0.979 73.551
W/o AdaIN-guided latent-shift 37.649 26.909 0.349 0.985 0.990 0.980 73.655
W/o AdaIN-guided attention-shift 39.589 27.506 0.388 0.985 0.988 0.983 72.958
W/o sliding-window consistent smoothing 38.626 26.907 0.384 0.982 0.989 0.970 73.701
W/ all components (ours) 37.152 26.134 0.369 0.986 0.991 0.980 74.014
DAVTG-Laion
W/o all components (baseline) 38.981 26.594 0.413 0.979 0.976 0.981 71.571
W/o localized latent blending 30.848 20.974 0.404 0.961 0.967 0.979 73.888
W/o AdaIN-guided latent-shift 31.001 21.766 0.362 0.985 0.989 0.980 73.987
W/o AdaIN-guided attention-shift 32.534 22.623 0.377 0.985 0.989 0.981 73.395
W/o sliding-window consistent smoothing 30.834 21.249 0.386 0.981 0.988 0.971 74.044
W/ all components (ours) 30.636 21.146 0.383 0.986 0.990 0.981 74.495
Moreover, we present additional visualizations of UniVST applied to four additional backbones, including SD-v2.1 [32], SD-v3.0 [43], SD-v3.5 [43], and Animatediff-v2 [44], to further highlight its generalization and scalability. As presented in Fig. 7, different backbones consistently generate similar stylization results, underscoring the stability of our framework. It is also worth noting that UniVST also achieves stable visual performance under the state-of-the-art rectified-flow models such as SD-v3.0 and SD-v3.5, in practice requiring only the replacement of the inversion technique, thereby further demonstrating its scalability.
Figure 8: Visual results for ablation studies. Various components in our framework have all played important roles.
5.4 Ablation Studies
In this section, we first validate the effectiveness of each core component in our proposed UniVST. Then, we conduct an in-depth analysis of hyper-parameter settings for the three modules: point-matching mask propagation, AdaIN-guided localized video stylization, and sliding-window consistency smoothing. All ablation experiments are conducted on SD-v1.5.
5.4.1 Ablation Study on Core Components
We have developed six distinct variants to assess the contribution of individual components in our proposed UniVST: (1) UniVST without all components (baseline), (2) UniVST without localized latent blending, (3) UniVST without AdaIN-guided latent-shift, (4) UniVST without AdaIN-guided attention-shift, (5) UniVST without sliding-window consistent smoothing and (6) UniVST with all components (ours). The quantitative outcomes are presented in Table III, while the qualitative results are displayed in Fig. 8. Our localized latent blending strategy effectively preserves the foreground style during the stylization process, demonstrating its efficacy quantitatively and qualitatively, without the need for additional models like SAM [35]. The AdaIN-guided latent-shift and attention-shift strategies are essential for style transfer. Notably, omitting the attention-shift strategy leads to incomplete transfer and blurred local details, as shown in Fig. 8. Moreover, our sliding-window consistency smoothing scheme enhances temporal consistency and effectively reduces video artifacts. Overall, the coordinated effect of these components yields the lowest ArtFID [17], indicating optimal content–style balance; the highest SSIM [55] and CLIP-I [56], reflecting effective local preservation; and the highest VBench score, showing superior overall quality across multiple dimensions, including consistency, aesthetics, art, and so on.
Figure 9: Comparison of stylization performance among our mask propagation strategy, SAM-H, and SAM-L on DAVTG-Laion dataset.
As shown in Fig. 3, our mask propagation strategy demonstrates spatial and temporal advantages over frame-by-frame segmentation with the SAM model [35]. To further assess the stylization performance of our mask propagation strategy, we perform local video style transfer using masks generated by three approaches: SAM-H, SAM-L, and our mask propagation. This ablation study is performed on the DAVTG-Laion dataset, and the corresponding quantitative comparison is presented in Fig. 9. Our mask propagation strategy outperforms SAM-L on the VBench metric, while achieving performance comparable to both SAM-L and SAM-H on the other three metrics, ArtFID, SSIM, and CLIP-I. These results highlight the robustness of our strategy in balancing content and style, preserving local stylization, and maintaining overall visual quality during the localized video style transfer process.
Figure 10: Quantitative results of mask propagation with different sampling ratios $r$ on DAVIS [48] dataset. Best viewed with zooming-in.
In the following section, we conduct an in-depth study of the three proposed modules on the DAVIS-Laion dataset to further substantiate their individual contributions, respectively.
5.4.2 Analysis of Point-Matching Mask Propagation
We first investigate the effect of down-sampling rates on propagation accuracy and inference time using the DAVIS dataset [48]. Detailed results are presented in Fig. 10. As the down-sampling rate $r$ decreases, the efficiency of our mask propagation improves significantly, while maintaining the accuracy of the final propagation results. Next, we also examine the impact of anchor frame selection on propagation accuracy and efficiency. As shown in Table IV, increasing the number of anchor frames improves propagation accuracy but also leads to longer propagation times. Moreover, we observe that as the number of frames increases, the accuracy gains diminish. After considering these trade-offs, we choose the first frame and the previous $n=9$ frames as the final hyper-parameter configuration. We also compare the results for different values of the hyper-parameter $k$ in $k$-NN [37]. As shown in Table V, the highest propagation accuracy is consistently achieved when $k=15$, without sacrificing the overall propagation efficiency.
TABLE IV: Accuracy and efficiency of mask propagation under different values of $n$ on DAVIS-Laion dataset, Bold fonts represent final configuration.
@@ 5.4.2 Analysis of Point-Matching Mask Propagation
n=1 0.5321 0.6505 2.365
@@
TABLE V: Accuracy and efficiency of mask propagation with different values of $k$ on DAVIS-Laion dataset. Bold fonts represent final configuration.
@@
TABLE VI: Comparison of our method with other training-free stylization methods on DAVIS-Laion dataset. Bold values indicate final configuration.
Methods K-V Repl. K-V AdaIN Ours ($\beta_{t}\uparrow$) Ours ($\beta_{t}\downarrow$)
[…]
5.4.3 Analysis of AdaIN-Guided Localized Video Stylization
We first investigate the effectiveness of our proposed AdaIN-guided attention-shift mechanism and the impact of the monotonicity of the time coefficient $\beta_{t}$. Specifically, we compare the stylized performance of four configurations: K-V replacement [17], K-V AdaIN, our method with increasing time coefficients, and our method with decreasing time coefficients. As shown in Table VI, using a decreasing time coefficient enables the AdaIN-guided attention-shift to outperform other training-free methods, such as K-V replacement, K-V AdaIN, or attention-shift with increasing time coefficients.
@@ 5.4.3 Analysis of AdaIN-Guided Localized Video Stylization
Timestep Intervals $\leq 0.6T$ $\leq 0.5T$ $\leq 0.4T$ $\leq 0.3T$ $\leq 0.2T$ $\geq\boldsymbol{0.4T}$ $\geq 0.5T$ $\geq 0.6T$ $\geq 0.7T$ $\geq 0.8T$
@@
Timestep Intervals $\boldsymbol{0.1T}\sim\boldsymbol{0.2T}$ $0.2T\sim 0.3T$ $0.4T\sim 0.5T$ $0.5T\sim 0.6T$ $0.7T\sim 0.8T$ $0.8T\sim 0.9T$
@@
Figure 11: Visual results of latent-shift and attention-shift with different timestep intervals on DAVIS-Laion dataset. Best viewed with zooming-in.
In addition, we further investigate the optimal timestep intervals for AdaIN-guided attention-shift and latent-shift. As shown in Table VII, applying attention-shift in the early and middle stages of denoising yields better results. In contrast, Table VIII shows that latent-shift performs better when applied in the late stages. Moreover, we present visual results of these ablation experiments, as shown in Fig. 11. For AdaIN-guided latent-shift, applying it in the early stages of denoising can cause issues such as content blurriness and semantic loss during the transfer process. In contrast, AdaIN-guided attention-shift produces better visual results when applied in the early stages of denoising than in the later stages.
Figure 12: Visual results of sliding-window consistent smoothing with different window sizes on DAVIS-Laion dataset. Best viewed with zooming-in.
@@ 5.4.4 Analysis of Sliding-Window Consistency Smoothing
Figure 13: Quantitative results of sliding-window consistent smoothing with varying timestep intervals. Best viewed with zooming-in.
@@ 5.5 Addtional Analysis
Bottleneck of Inference Time. In Table I, our proposed UniVST exhibits the longest inference time among all compared methods, revealing a key computational bottleneck. To better understand this limitation, we perform a detailed breakdown of the inference time distribution across the core modules of UniVST, as illustrated in Fig. 14. These modules include: DDIM Inversion, Point-Matching Mask Propagation, AdaIN-Guided Localized Video Stylization, and Sliding-Window Consistent Smoothing. Our analysis indicates that the sliding-window consistent smoothing module accounts for about 37.1% of the total overhead, primarily due to the dynamic estimation of inter-frame optical flow within each sliding window during specific denoising intervals. This component involves a trade-off: although it enhances temporal consistency in a training-free manner during the denoising process, the frequent estimation of optical flow inevitably introduces additional computational overhead. To address the above trade-off, we propose two training-free alternatives: (1) employing a 3D VAE decoder to reconstruct temporally coherent video from the denoised latent representation; and (2) incorporating a plug-and-play pretrained adapter after the self-attention module for temporal modeling, similar to the motion module in AnimateDiff [44].
Constraints in Mask Propagation. We present qualitative examples of our mask propagation results in Fig. 15. In the short-term propagation setting, we evaluate scenarios involving fast-moving foreground objects, deformations, and occlusions. Despite these complexities, our method demonstrates precise object tracking and robust mask alignment, consistently preserving accuracy across diverse conditions. For long-term propagation, our approach continues to deliver strong performance. The integration of anchor frames enables stable and accurate mask propagation over extended temporal distances, mitigating typical degradation issues associated with long sequences. Although our point-matching mask propagation strategy enables zero-shot object localization, it presents several limitations: 1. Challenges in Multi-Object Scenarios. In scenes containing multiple similar objects (e.g., two individuals), the semantic similarity between instances can lead to confusion during propagation. This often results in incorrect mask associations, particularly as the propagation distance increases, causing error accumulation over time. While our current evaluation focuses on diverse yet single-object scenes—where such issues have minimal impact—extending the method to robustly handle multi-object scenarios remains an important direction for future work. 2. Limitations in Fine-Grained Mask Propagation. Although the propagation achieves an average Dice score of approximately $0.87$, it occasionally fails to capture fine object boundaries, leading to incomplete or imprecise tracking. As shown in Fig. 15, while the method effectively handles fast motion and occlusions, it struggles to retain detailed edge information. Nonetheless, our localized latent blending strategy, which performs dynamic region-wise fusion during the denoising phase, effectively mitigates this limitation by reducing reliance on high-precision masks.